OpenStudy (anonymous):

determine whether the equation is an identity cos^4θ+1-sin^4θ=2cos^3θ

OpenStudy (anonymous):

\[is the equation is \cos4\theta+1-\sin4\theta=2\cos3\]

OpenStudy (anonymous):

\[\cos ^4 \theta+1-\sin ^4\theta=2 \cos ^3\theta\]

OpenStudy (anonymous):

\[x^4-y^4=(x^2-y^2)(x^2+y^2)\] \[\cos ^4\theta-\sin ^4\theta=(\cos ^2\theta-\sin ^2\theta)(\cos ^2\theta+\sin ^2\theta)=(\cos ^2\theta-\sin ^2\theta)(1)\] \[\cos ^2\theta-\sin ^2\theta+1=\cos ^2\theta+1-\sin ^2\theta=\cos ^2\theta+\cos^2\theta=2\cos^2 \theta\]

OpenStudy (anonymous):

so no